Using the data in the following table, and the fact that the correlation of A and B is 0.48, calculate the volatility standard deviation of a portfolio that is 70% invested in stock A and 30% invested in stock B Realized Returns Year 2008 2009 2010 2011 2012 2013 Stock A -10% 20% 5% -5% 2% 996 Stock B 21% 30% 7% -3% -8% 25% The standard deviation of the portfolio is 96. (Round to two decimal places.)
Answer #1

SEE THE SCREENSHOT. ANY DOUBTS, HAPPY TO HELP YOU. THANK YOU. THUMBS UP PLEASE.

standard deviation is calculated using excel function : stdev.s.

YEAR STOCK A STOCK B 2008 2009 2010 2011 2012 2013 21 0.7 0.3 0.48 10.5972 15.6461 -10 20 WA- 30 rAB = sd(A) = sd(B)- -5 -3 9 25 STANDARD DEVIATION 10.5972 15.6461 PORTFOLIO SD sqrt [ (wA)A2*(sd(A))A2 + (wB)^2*(sd(B))A2 + 2 wA*wB*(sd(A))(*sd(B)) rAB] PORTFOLIO SD SQRT [(0.7)A2 (10.5972)A2 (0.3)A2*(15.6461)A2 2*(0.7) (0.3)*(10.5972) (15.6461)* (0.48)] PORTFOLIO SD - 10.51